OpenStudy (anonymous):

A sled is pushed along an ice covered lake. It has some initial velocity before coming to a rest in 15 m. It took 23 seconds before the sled and rider come to a rest. If the rider and sled have a combined mass of 52.5 kg, what is the magnitude (amount) and direction of stopping force? OK, so I have no idea what to do tbh. HELP PLEASE

OpenStudy (anonymous):

@pooja195 ?

OpenStudy (anonymous):

The most difficult part for me is the initial velocity. I don't know what it is. I don't know how to find it.

pooja195 (pooja195):

\[\huge~\rm~u=v-at\] u = Initial velocity, v = Final Velocity, t = time taken, a = acceleration. This is the formula for intial velocity but in order to solve this we need the other info...

OpenStudy (anonymous):

final velocity is 0. time taken is 23 minutes. acceleration is idk. but it is negative

pooja195 (pooja195):

Wouldn't it be -15?

OpenStudy (anonymous):

but it took 23 minutes right?

pooja195 (pooja195):

Its decelerating... It has some initial velocity before `coming to a rest in 15 m`.

pooja195 (pooja195):

Im not sure :/

OpenStudy (anonymous):

m = meters

OpenStudy (irishboy123):

you can work this from the equations of motion do you know them?

jimthompson5910 (jim_thompson5910):

http://www.cliffsnotes.com/~/media/387da48a5492469fb38fd6dfeb115d4e.ashx?la=en we don't know the acceleration, so you use the equation with an X in the acceleration "a" column that is equation 2 which says \[\Large d=\frac{1}{2}(v_0+v_f)*t\] d = 15 meters V0 = unknown (but can be solved for) Vf = 0 meters per second t = 23 seconds

OpenStudy (anonymous):

yes. I know acceleration, which idk if i can write on here directly, f = ma, and v = d/t and magnitude is an arrow, and i need the # and the direction is the opposite i think.

OpenStudy (irishboy123):

jt has it !

OpenStudy (anonymous):

so 15 = 1/2 ^ ( [unknown] + 0) ^ 23?

jimthompson5910 (jim_thompson5910):

Once you know V0, you can find the acceleration Then you'll use the acceleration to compute F in `F = m*a` to find the stopping force (aka frictional force)

OpenStudy (anonymous):

I think it is 1.3 m/s because i divided 15 by 2, and then by 23 and got 1.3. Is that right?

OpenStudy (anonymous):

so it is 0-1.3 / 23-0?

OpenStudy (anonymous):

no that doesnt sound right

jimthompson5910 (jim_thompson5910):

why not? you're using \[\Large a = \frac{\Delta v}{\Delta t}\] which is correct

OpenStudy (anonymous):

but @pooja195 got the answer, and it was like 2.8 N

jimthompson5910 (jim_thompson5910):

you should find that \[\Large a = \frac{\Delta v}{\Delta t} \] \[\Large a = \frac{v_f-v_0}{\Delta t}\] \[\Large a = \frac{0-1.304347826}{23}\] \[\Large a = -0.05671077504\] So, F = m*a F = 52.5*(-0.05671077504) F = -2.9773156896

OpenStudy (anonymous):

OMG YOU ARE LITERALLY THE BEST OMGOMGOMGOMGOGOGOMGMGMOGMOMGOMGOOGMGMOGMOGOMGOGM <3

jimthompson5910 (jim_thompson5910):

The magnitude is simply the absolute value of the force The direction of the stopping force is the complete opposite of the motion of the sled

OpenStudy (anonymous):

so like ------> -2.97 N The direction is because it is negative.

jimthompson5910 (jim_thompson5910):

`but @pooja195 got the answer, and it was like 2.8 N` I'm guessing @pooja195 rounded somewhere to get 2.8

jimthompson5910 (jim_thompson5910):

I don't know what you mean by that arrow. It doesn't say which direction this sled is moving (left or right)

OpenStudy (anonymous):

So technically, it is right, and wrong?

jimthompson5910 (jim_thompson5910):

If the sled is moving left, then the stopping force is pointing to the right If the sled is moving right, then the stopping force is pointing to the left

jimthompson5910 (jim_thompson5910):

I would just say the stopping force is opposite the direction of motion (until you can get more info about where the sled is going)